If 3 is added to twice a number and this sum is multiplied by 2, the result is the same as if the number is multiplied by 0 and 2 is added to the product. What is the number?
Answer by sudhanshu_kmr(1152)


(3+ 2*x )* 2 = x* 0 + 2
=> 6 + 4x = 2
=> 4x = 2-6
=> x = -4/4 = -1

number is: -1